A protein sample is analyzed by a kjeldahl process during which ammonia is formed. Ammonia
formed is titrated coulometrically with electrogenerated hypobromite (OBr). If the titration is
performed using 19.3 milliamperes current and end point occur at 2 minutes. Calculate the
milligrams of protein present in the sample if one gram of protein nitrogen is contained in each
6.25gof protein (H=1 and N=14).
1
Expert's answer
2020-09-24T11:21:36-0400

Equivalent mass of NH3 17 g (NH3 is monobasic)

"m = ZIT"

for NH3 , "z= \frac{17g}{faradey}"


"= \frac{17g}{96500c}"


Mass deposited "=" "\frac{17g}{96500c}\times 19.3\times10" -3"A\times2\times60" sec


"= 0.408\times10" -3g


In NH3 17 g contains 3g of H and 15 g of N


17g NH3 -------------- 15g N


1g NH3 ----------------"\frac{15}{7}N"


NH3 ---------------- 0.36"\times 10" -3N


0.36"\times 10" -3N contain in 6.25"\times" 0.36"\times 10" -3g protein

"= 6.25\times0.36" mg protein

"= 2.25 mg" protein